OpenStudy (anonymous):

I need this answered quick find three consecutive odd interageres such that twice the sum of the smaller two interagers is 25 more than 3 times the largest

OpenStudy (anonymous):

\[2(x+x+2)=3(x+4)+25\] maybe that is better

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

go with the second one you get \[4x+4=3x+37\] if my math is correct

OpenStudy (anonymous):

Ok thanks